OpenStudy (anonymous):

If I graphed (6, 9) and (-2, -3) on a coordinate plane, how do I find the slope of the line, and what is the relationship between the two? Please explain how I can solve this.

OpenStudy (anonymous):

slope is just (y2-y1)/(x2-x1)

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

-3-9 = -12 and -2-6 = -8 -12/-8 = 3/2

OpenStudy (anonymous):

Let's say I put (6, 9) as points on a graph, if I wrote a line between those two points, what would be the point of that specific line? That's what I really wanted to know.

OpenStudy (anonymous):

which two points

OpenStudy (anonymous):

This

OpenStudy (anonymous):

you can only find the slope of that line

OpenStudy (anonymous):

Yes, how do I get the slope of that line?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

i'll take a pic of the work

OpenStudy (anonymous):

That's strange because on my work here, it asks for the slope of (6,9) and (-2, 3), yet you're combining both of them.

OpenStudy (anonymous):

Slope of the line, that is.